ne 1: 80 81 82 78 83







Mine 2: 77 76 78 80 76 78







Use 0.05 level of significance to test whether the difference between the means of these two samples is significant.

1
Expert's answer
2022-01-14T09:14:28-0500

X = Represent mine 1

Y = Represent mine 2

"Mean (\bar{X})=\frac{\Sigma X}{n_1}=\frac{404}{5}=80.8"


"Mean (\bar{Y})=\frac{\Sigma Y}{n_2}=\frac{465}{6}=77.5"


Sample standard deviation for each sample:

"S_X=\sqrt{\frac{(80-80.8)^2+(81-80.8)^2+(82-80.8)^2+(78-80.8)^2+(83-80.8)^2}{5-1}} =1.924"


"S_Y=\sqrt{\frac{(77-77.5)^2+(76-77.5)^2+(78-77.5)^2+(80-77.5)^2+(76-77.5)^2+(78-77.5)^2}{6-1}} =1.517"


Following steps are taken to test whether the difference between the means of these two samples is significant at  0.05 level of significance:

Null & Alternative Hypothesis:

"H_0:\mu_1=\mu_2"

"H_a:\mu_1\not =\mu_2"

Test Statistic:

Sample size is less than 30 for each of two samples. So, t-test for two-samples is used to test hypothesis.

"t=\frac{\bar{X}-\bar{Y}}{\sqrt{\frac{S_X}{n_1}+\frac{S_Y}{n_2}}}"


"t=\frac{80.8-77.5}{\sqrt{\frac{1.924}{5}+\frac{1.517}{6}}}=4.133"

Critical Value:

"df=n_1+n_2-2=5+6-2=9"

Critical value at 0.05 level of significance and degree of freedom is 9 would be:

"t_c=1.833"

Decision Making:

Hence, test statistic is greater than critical value, so the null hypothesis is rejected.

Conclusion:

Based on this, it can conclude that the difference between the means of these two samples is significant.